OpenStudy (anonymous):

find dy/dx if lnx-lny^2+sinxy=8

OpenStudy (amistre64):

what is the derivative of ln(x) ?

OpenStudy (amistre64):

and then of: -lny^2 and of sin(xy) and lastly of 8

OpenStudy (anonymous):

can i send the wolfram ans

OpenStudy (anonymous):

for this question

OpenStudy (lgbasallote):

why?

OpenStudy (anonymous):

for clearifying

OpenStudy (anonymous):

I had trouble with the last term, but found an answer here: http://www.scienceforums.net/topic/23152-implicit-diff/

OpenStudy (amistre64):

the last term? really? 8' = 0

OpenStudy (anonymous):

second to last :)

OpenStudy (lgbasallote):

clarifying what? there wasnt even anything that needed to be clarified yet...

OpenStudy (anonymous):

this question but i think the ans is wrong in wolfram

OpenStudy (amistre64):

sin(xy); use chain rule cos(xy) * (xy)' ; use product rule cos(xy) * (x'y + xy') ; since dx/dx=x' = 1, ignore the x's cos(xy)(y+xy')

OpenStudy (anonymous):

@ParthKohli

OpenStudy (lgbasallote):

usually askers appreciate being guided and taught rather than being given an answer @best.shakir :)

OpenStudy (anonymous):

i also have doubt in this question

OpenStudy (anonymous):

@amistre64 thank u

OpenStudy (amistre64):

yalls welcome :)

OpenStudy (anonymous):

just to add little bit here this requires Implicit differentiation. and for the second term derivative should be \[\Large \frac{d}{dx}\ln(y^2) =\frac{1}{y^2}*2y \frac{dy}{dx}\].